# xLSTM Scaling Laws: Competitive Performance with Linear Time-Complexity

### v1

@@ Abstract
Scaling laws play a central role in the success of Large Language Models (LLMs), enabling the prediction of model performance relative to compute budgets prior to training. While Transformers have been the dominant architecture, recent alternatives such as xLSTM offer linear complexity with respect to context length while remaining competitive in the billion-parameter regime. We conduct a comparative investigation on the scaling behavior of Transformers and xLSTM along the following lines, providing insights to guide future model design and deployment. First, we study the scaling behavior for xLSTM in compute-optimal and over-training regimes using both IsoFLOP and parametric fit approaches on a wide range of model sizes (80M-7B) and number of training tokens (2B-2T). Second, we examine the dependence of optimal model sizes on context length, a pivotal aspect that was largely ignored in previous work. Finally, we analyze inference-time scaling characteristics. Our findings reveal that in typical LLM training and inference scenarios, xLSTM scales favorably compared to Transformers. Importantly, xLSTM’s advantage widens as training and inference contexts grow.
Our code and data is available at: github.com/NX-AI/xlstm_scaling_laws
@@ 1 Introduction
Recent works (Sardana et al., 2024; Gadre et al., 2024) rightfully argue that these scaling laws are nevertheless limited by their neglect of inference costs. Consequently, these works focus on performance investigations on models that are trained in the so-called over-training regime, i.e. on more tokens than would be optimal in terms of pre-taining compute. Importantly, these works and subsequent ones focus on Transformer architectures (Vaswani et al., 2017). In these architectures, the attention mechanism inflicts computational costs during training and inference that are quadratic in terms of context length. Besides the associated economic and ecological costs, this quadratic scaling is prohibitive for a large range of application areas in which models are deployed on devices with limitations on available memory, energy, or allowable TFTT. Even on GPUs that are dedicated to LLMs this scaling property of Transformers represents a limitation in task that require very long contexts, like reasoning (Muennighoff et al., 2025). Consequently, the development of LLM architectures that mitigate the attention mechanism is an active area of research (Gu & Dao, 2024; Beck et al., 2024; Lieber et al., 2024). While these architectures were demonstrated to be scalable into the billion-parameter regime (Zuo et al., 2024; Beck et al., 2025b), there is so far no systematic comparison between linear complexity LLM architectures, i.e. LLMs that scale linearly in computational costs w.r.t. context lengths, and transformer-based LLMs with quadratic complexity.
@@
Our investigation shows, that xLSTM models are Pareto-dominant in terms of cross-entropy loss over Transformer models (Fig. 1), enabling models that are both better and cheaper. We find that, for a given training compute budget, compute-optimal xLSTM models are larger (Fig. 4), i.e. have more parameters, than compute-optimal Transformer models. During inference, xLSTMs are faster than same-sized Transformers (Fig. 6), and their performance advantage grows with context length due to Transformers’ quadratic time complexity.
@@ 2.1 Background on Scaling Laws
Previous works on transformer scaling laws commonly approximate compute costs with $C(N,D)=6ND$ FLOPs (Kaplan et al., 2020; Hoffmann et al., 2022; Gadre et al., 2024; Sardana et al., 2024). This approximation ignores the FLOPs associated to the attention mechanism and covers only the feed-forward network contributions. Recently, several works (DeepSeek-AI, 2024a; Busbridge et al., 2025; Li et al., 2025) pointed out that this approximation is not justified for sufficiently large context lengths and models. For the purpose of this work, this approximation is even less suitable since it neglects entirely the difference between linear and quadratic time-complexity models. Hence, we adopt a more precise calculation of $C(N,D)$ as provided in Appendix C.3 that accurately captures the differences in computational complexity between model classes.
@@ 3.1 Experimental Setup
For both model classes we use the same training recipe derived from the xLSTM 7B training recipe (Beck et al., 2025b); details are given in Appendix A.1. The overall number of training steps is determined by the FLOP budget or token-to-parameter ratio of the specific experiment. As training dataset, we use DCLM-Baseline, a collection of high-quality filtered web documents (Li et al., 2024), tokenized with the GPT-NeoX tokenizer (Black et al., 2022) into sequences of length 8192, unless specified otherwise.
@@
Using the above defined architecture and training recipe, we produce a large dataset of training runs for our scaling law study totaling 672 individual runs (292 for Llama, 380 for xLSTM). The dataset contains model sizes ranging from 80M to 7B parameters trained with compute budgets ranging from $2.8\times 10^{18}$ to $8.5\times 10^{22}$ FLOPs on 2B to 2T tokens. This amounts to a total compute budget spent for this dataset of $3.2\times 10^{23}$ FLOPs. Our dataset is divided in into runs from two different training configurations: IsoFLOP and Token/Param. For the IsoFLOP configuration, we vary model parameters and training tokens subject to fixed compute budgets for three different context lengths. In the Token/Param configuration, we vary the number of training tokens for a set of fixed model sizes. We show our dataset as $\{N,D,C\}$ points in Figure 2. xLSTM’s linear scaling preserves training tokens with longer contexts (overlapping IsoFLOP lines), whereas Transformer’s quadratic scaling reduces them.
@@
We begin our study with the question: Given a fixed training compute budget, which model architecture performs better (in terms of cross-entropy loss)? To answer this question, we define a grid of model and dataset sizes with pre-defined token-to-parameter ratios of $[22,44,110,220,550,1100,2200]$ and train Transformer and xLSTM models for each point in the grid. This forms the Token/Param subset in our dataset of training runs (see Sec. 3.1). We then use our FLOP calculations in Appendix C.3 and plot validation loss over FLOPs in a log-log plot in Figure 1.
@@ 3.4 Compute-Optimal xLSTM Models are Larger
Analogous results are shown in Figure 9 in the appendix for the number of training tokens of compute-optimal models. We find that compute-optimal xLSTM and Transformer models are trained on a similar number of training tokens $\hat{D}^{*}(H)$.
@@
The compute-optimal models in Figure 4 (left) fall close to a single shared line for the Transformer and xLSTM models. This suggests that for compute-optimal models, there is a universal relationship between performance and model size for xLSTM and Transformer models. From this perspective, the fact that compute-optimal xLSTM models are larger for a given compute budget can be regarded as a heuristic explanation for the superior performance of xLSTM.
Figure 4: Varying model size and tokens with a fixed compute budget (IsoFLOP). Left: IsoFLOP profiles for varying number of model parameters with a marker at the minimum $N^{*}$ of the fitted polynomial. Right: Power-law fit $N^{*}(H)=A^{\prime}\cdot H^{a}$ for the compute optimal number of model parameters. Our setup reproduces the power-law exponent $a$ for Transformers in Porian et al. (2024).
@@ 3.5 Compute-optimal xLSTM model size remains stable across Context Lengths
The main difference between the model architectures in this study is their scaling in FLOPs with context length: Transformers scale quadratically, due to the self-attention, while xLSTMs scale linearly. This implies that, in Transformers, an increasing fraction of compute is devoted to attention as sequence length grows, whereas in xLSTMs the recurrent updates consume only a modest portion of the total compute. In this section, we investigate, therefore, the impact of the context length on compute-optimal model and dataset sizes. We add experiments with context lengths 2048 and 16384 in the IsoFLOP training configuration and then fit the power-laws to each context length for both models, analogously to Section 3.4. We note that the losses are not comparable across different context lengths since the corresponding datasets are not identically distributed.
@@
In Figure 5 we show the IsoFLOP profiles for varying model sizes and three different context lengths and compute budgets, including their power-law fits $\hat{N}^{*}(H)$ in the rightmost plot. We observe that with increasing context lengths the compute-optimal model size of Transformers drops significantly, while for xLSTM it drops only mildly. These results suggest that for Transformers, a growing fraction of compute is consumed by attention operations as sequence length increases, whereas in xLSTMs most FLOPs remain allocated to depth and hidden dimensions. As context length grows, compute-optimal Transformers exhibit a significantly larger performance drop than xLSTM. In Figure 10 in Appendix A.5 we show the corresponding IsoFLOP profiles and power-law fits $\hat{D}^{*}(H)$ for the optimal number of training tokens. We observe similar trends as for the model size: The compute-optimal number of training tokens decreases markedly with larger context length for Transformer models and for xLSTM it slightly increases.
Figure 5: Left: IsoFLOP curves as a function of model parameters at 3 different context lengths. Right: Plot of the power-law fits for the compute optimal number of parameters dependent on the compute budget $N^{*}(H)$. Colors indicate compute budget and marker types indicate the model types.
@@ 4 Inference Scaling Behavior
The scaling laws analysis in Section 3 is motivated by the goal of the optimal design of pre-training runs for LLMs. However, these considerations neglect inference efficiency. When deploying LLMs at large scale, inference costs and performance are critical aspects. Hence Pope et al. (2023) investigate the inference efficiency of transformer-based LLMs in terms of three criteria: compute, latency, and throughput. More recently Sardana et al. (2024) provided a scaling law analysis of Transformers that extend the pre-training compute optimality consideration (Eq. (1)) to also account for inference compute. This work presents an even more comprehensive analysis in terms of the attainable latency, i.e., time to first token, and the step time during generation. We complement our empirical findings with a quantitative model of a lower bound on time to first token and step time, using the detailed calculation of FLOPs (App. C.3) and MemOps (App. C.4) for both model architectures.
@@
For the prefill stage, the key performance metric is the time to first token (TTFT). Prefill speed is primarily determined by how well the model can maintain a low TTFT while handling large batch sizes and long input sequences. During the generation stage, the key performance metric is the step time, i.e. how long it takes to obtain the next token given the current (potentially batched) sequence. For Transformers, the quadratic complexity of the attention mechanism with respect to the prefill length (App. C.3.3) implies that TTFT is expected to scale quadratically in terms of the prefill length. In terms of step time we expect linear scaling with respect to the prefill length, as each decoding step involves attention over the entire KV cache. For xLSTMs, in contrast, we expect linear scaling of TTFT and step time that is independent of the prefill length.
@@ 4.2 Modeling Inference Runtimes
In our analysis, the inference processes are characterized by the associated number of floating point operations $\text{FLOPs}_{\text{algo}}$ and the number of memory operations $\text{Bytes}_{\text{mem,algo}}$ measured in bytes that are read or written. We provide calculations of these two quantities for xLSTM and for Transformers in Appendix C. Importantly, these calculations capture the difference between xLSTM and Transformers in the dependence of $\text{FLOPs}_{\text{algo}}$ and $\text{Bytes}_{\text{mem,algo}}$ on the context length $T$. Based on these calculated quantities, we model the runtimes associated with the floating point and memory operations as:
@@
where $\alpha_{\text{ eff}{{}}}$ is the effective rate of FLOPs/s, $\beta_{\text{ eff}{{}}}$ is the effective rate of Bytes/s, and $\epsilon$ is a constant overhead when running the inference processes on the GPU. Depending on the model type, model size, prefill length, batch size and inference stage (prefill or generate), either $\tau_{\text{ FLOPs}{{\text{,algo}}}}$ or $\tau_{\text{ mem}{{\text{,algo}}}}$ is the dominant contributor to the runtime. We outline in Appendix D.1 how this is determined based on the roofline model. Using empirical runtime measurements, we then fit one of the two models depending on which one is expected to yield the dominant runtime contribution. Each fit corresponds to a specific model type, size, and inference stage, and is evaluated over varying batch sizes and prefill lengths. As evidenced by the fits to empirical TTFT (App. D.2) and step time measurements (App. D.3), our model provides an accurate description of the observed inference runtimes for both architectures and explains the empirically observed runtimes in Figure 6.
@@ 5 Related Work
The empirical scaling behavior of Deep Learning models w.r.t the size of their model parameters and training data has been actively researched (Hestness et al., 2017; Rosenfeld et al., 2020; Henighan et al., 2020; Alabdulmohsin et al., 2022; Caballero et al., 2023). Such scaling laws have been demonstrated across many tasks and data modalities (Tan & Le, 2019; Ghorbani et al., 2022; Zhai et al., 2022; Abnar et al., 2022; Ardalani et al., 2022; Gao et al., 2023) However, beginning with Kaplan et al. (2020) and Hoffmann et al. (2022), the main objective has been guidance on how to optimally scale Large Language Models with Transformers. Follow-up work investigated the data constrained setting (Muennighoff et al., 2023), the effect of data pruning (Sorscher et al., 2022), extreme token per parameter ratios (Gadre et al., 2024). Furthermore, replication efforts regarding the scaling laws established in Kaplan et al. (2020) and Hoffmann et al. (2022) have been performed in order to reconcile their findings (Besiroglu et al., 2024; Pearce & Song, 2024; Porian et al., 2024). Critical practical considerations such as specific architectures and hyperparameters on the resulting scaling laws have been investigated (McLeish et al., 2025). The recent survey Li et al. (2025) gives a comprehensive overview and give practical guidelines in establishing scaling laws. Scaling laws have also been investigated theoretically, providing justification for the functional forms used in practice (Amari et al., 1992; Amari, 1993; Seung et al., 1992; Amari & Murata, 1993; Cortes et al., 1993; Yarotsky, 2018; Liang et al., 2020; Sharma & Kaplan, 2022; Hutter, 2021; Bahri et al., 2024).

### v2

@@ Abstract
Scaling laws play a central role in the success of Large Language Models (LLMs), enabling the prediction of model performance relative to compute budgets prior to training. While Transformers have been the dominant architecture, recent alternatives such as xLSTM offer linear complexity with respect to context length while remaining competitive in the billion-parameter regime. We conduct a comparative investigation on the scaling behavior of Transformers and xLSTM along the following lines, providing insights to guide future model design and deployment. First, we study the scaling behavior for xLSTM in compute-optimal and over-training regimes using both IsoFLOP and parametric fit approaches on a wide range of model sizes (80M-7B) and number of training tokens (2B-2T).  [^1] The code and data to reproduce our analyzes and figures is available at:
https://github.com/NX-AI/xlstm_scaling_laws
Second, we examine the dependence of optimal model sizes on context length, a pivotal aspect that was largely ignored in previous work. Finally, we analyze inference-time scaling characteristics. Our findings reveal that in typical LLM training and inference scenarios, xLSTM scales favorably compared to Transformers. Notably, xLSTM models consistently Pareto-dominate Transformer models, delivering lower cross-entropy loss for the same compute budget.
@@ 1 Introduction
Recent works (Sardana et al., 2024; Gadre et al., 2024) rightfully argue that these scaling laws are nevertheless limited by their neglect of inference costs. Consequently, these works focus on performance investigations on models that are trained in the so-called over-training regime, i.e. on more tokens than would be optimal in terms of pre-taining compute. Importantly, these works and subsequent ones focus on Transformer architectures (Vaswani et al., 2017). In these architectures, the attention mechanism inflicts computational costs during training and inference that are quadratic in terms of context length. Besides the associated economic and ecological costs, this quadratic scaling is prohibitive for a large range of application areas in which models are deployed on devices with limitations on available memory, energy, or allowable TFTT. Even on GPUs that are dedicated to LLMs this scaling property of Transformers represents a limitation in task that require very long contexts, like reasoning (Muennighoff et al., 2025). Consequently, the development of LLM architectures that mitigate the attention mechanism is an active area of research (Gu and Dao, 2024; Beck et al., 2024; Lieber et al., 2024). While these architectures were demonstrated to be scalable into the billion-parameter regime (Zuo et al., 2024; Beck et al., 2025b), there is so far no systematic comparison between linear complexity LLM architectures, i.e. LLMs that scale linearly in computational costs with respect to context lengths, and transformer-based LLMs with quadratic complexity.
@@
Our investigation shows, that xLSTM models Pareto-dominate Transformer models in the compute–loss trade-off (Fig. 1), enabling models that are both better and cheaper. We find that, for a given training compute budget, compute-optimal xLSTM models are larger (Fig. 4), i.e. have more parameters, than compute-optimal Transformer models. During inference, xLSTMs are faster than same-sized Transformers (Fig. 6), and their performance advantage grows with context length due to Transformers’ quadratic time complexity.
@@ 2.1 Background on Scaling Laws
Previous works on transformer scaling laws commonly approximate compute costs with $C(N,D)=6ND$ FLOPs (Kaplan et al., 2020; Hoffmann et al., 2022; Gadre et al., 2024; Sardana et al., 2024). This approximation ignores the FLOPs associated to the attention mechanism and covers only the feed-forward network contributions. Recently, several works (DeepSeek-AI, 2024a; Busbridge et al., 2025; Li et al., 2025) pointed out that this approximation is not justified for sufficiently large context lengths and models. For the purpose of this work, this approximation is even less suitable since it neglects entirely the difference between linear and quadratic time-complexity models. Hence, we adopt a more precise calculation of $C(N,D)$ as provided in Appendix B.3 that accurately captures the differences in computational complexity between model classes.
@@ 3.1 Experimental Setup
For both model classes we use the same training recipe derived from the xLSTM 7B training recipe (Beck et al., 2025b). The recipe uses the AdamW optimizer ($\beta_{1}=0.99$, $\beta_{2}=0.95$, $\epsilon=10^{-8}$), weight decay $0.1$ and gradient clipping norm $0.5$. The learning rate scheduler has three stages, linear warm-up, cosine decay to $10\%$ of the peak learning rate, and linear cool-down. For varying compute budgets, we scale the steps in the second stage while the first and third remain fixed. Further details are given in Appendix A.1. The overall number of training steps is determined by the FLOP budget or token-to-parameter ratio of the specific experiment. As training dataset, we use DCLM-Baseline, a collection of high-quality filtered web documents (Li et al., 2024), tokenized with the GPT-NeoX tokenizer (Black et al., 2022) into sequences of length 8192, unless specified otherwise. We use grain [^2] https://google-grain.readthedocs.io ( FirstFitPackIterDataset) to prepare batches with sequence packing, particularly first-fit packing, which avoids splitting, but adds padding tokens.
@@
Using the above defined architecture and training recipe, we produce a large dataset of training runs for our scaling law study totaling 672 individual runs (292 for Llama, 380 for xLSTM). The dataset contains model sizes ranging from 80M to 7B parameters trained with compute budgets ranging from $2.8\times 10^{18}$ to $8.5\times 10^{22}$ FLOPs on 2B to 2T tokens. This amounts to a total compute budget spent for this dataset of $3.2\times 10^{23}$ FLOPs. Our dataset is divided in into runs from two different training configurations: IsoFLOP and Token/Param. For the IsoFLOP configuration, we vary model parameters and training tokens subject to fixed compute budgets for three different context lengths. In the Token/Param configuration, we vary the number of training tokens for a set of fixed model sizes. We show our dataset as $\{N,D,C\}$ points in Figure 2. xLSTM’s linear scaling preserves training tokens with longer contexts (overlapping IsoFLOP points), whereas Transformer’s quadratic scaling reduces them.
@@
We begin our study with the question: Given a fixed training compute budget, which model architecture performs better (in terms of cross-entropy loss)? To answer this question, we define a grid of model and dataset sizes with pre-defined token-to-parameter ratios of $[22,44,110,220,550,1100,2200]$ and train Transformer and xLSTM models for each point in the grid. This forms the Token/Param subset in our dataset of training runs (see Sec. 3.1). We then use our FLOP calculations in Appendix B.3 and plot validation loss over FLOPs in a log-log plot in Figure 1.
@@ 3.4 Compute-Optimal xLSTM Models are Larger
Figure 4: Varying model size and tokens with a fixed compute budget (IsoFLOP). Left: IsoFLOP profiles for varying number of model parameters with a marker at the minimum $N^{*}$ of the fitted polynomial. Right: Power-law fit $N^{*}(H)=A^{\prime}\cdot H^{a}$ for the compute optimal number of model parameters. Our setup reproduces the power-law exponent $a$ for Transformers established in Porian et al. (2024). The compute-optimal model size of xLSTMs is larger than for Transformers.
@@
Analogous results are shown in Figure 9 in the appendix for the number of training tokens of compute-optimal models. We find that compute-optimal xLSTM and Transformer models are trained on a similar number of training tokens $\hat{D}^{*}(H)$. In Appendix E, we show the estimated optimal training FLOPs and training tokens for various model sizes.
@@
The compute-optimal models in Figure 4 (left) fall close to a single shared line for the Transformer and xLSTM models. This suggests that for compute-optimal models, there is a universal relationship between performance and model size for xLSTM and Transformer models. From this perspective, the fact that compute-optimal xLSTM models are larger for a given compute budget can be regarded as a heuristic explanation for the superior performance of xLSTM. The reason why xLSTMs can be larger is the reduced computational complexity of their recurrent sequence-mixing operation compared to the self-attention operation in Transformers. As this main operation is cheaper, more compute can be allocated to the rest of the model, e.g. increased number of layers or embedding dimension.
@@ 3.5 Compute-optimal xLSTM model size remains stable across Context Lengths
The main difference between the model architectures in this study is their scaling in FLOPs with context length: Transformers scale quadratically, due to the self-attention, while xLSTMs scale linearly. This implies that, in Transformers, an increasing fraction of compute is devoted to attention as sequence length grows, whereas in xLSTMs the recurrent updates consume only a modest portion of the total compute. In this section, we investigate, therefore, the impact of the context length on compute-optimal model and dataset sizes. We add experiments with context lengths 2048 and 16384 in the IsoFLOP training configuration and then fit the power-laws to each context length for both models, analogously to Section 3.4. We note that the losses are not directly comparable across different context lengths since we use sequence packing for the construction of our training and validation datasets. Hence, for larger context lengths, longer documents can be packed into a batch, effectively changing the data distribution.
@@
In Figure 5 we show the IsoFLOP profiles for varying model sizes and three different context lengths and compute budgets, including their power-law fits $\hat{N}^{*}(H)$ in the rightmost plot. We observe that with increasing context lengths the compute-optimal model size of Transformers drops significantly, while for xLSTM it drops only mildly. These results suggest that for Transformers, a growing fraction of compute is consumed by attention operations as sequence length increases, whereas in xLSTMs most FLOPs remain allocated to depth and hidden dimensions. In Figure 10 in Appendix A.5 we show the corresponding IsoFLOP profiles and power-law fits $\hat{D}^{*}(H)$ for the optimal number of training tokens. We observe similar trends as for the model size: The compute-optimal number of training tokens decreases markedly with larger context length for Transformer models and for xLSTM it slightly increases.
Figure 5: Left: IsoFLOP curves as a function of model parameters at 3 different context lengths. Right: Plot of the power-law fits for the compute optimal number of parameters dependent on the compute budget $N^{*}(H)$. Colors indicate compute budget and marker types indicate the model types. The compute optimal model size for Transformers gets smaller for larger context lengths, while the compute optimal model size for xLSTM remains similar across context lengths.
@@ 4 Inference Scaling Behavior
The scaling laws analysis in Section 3 is motivated by the goal of the optimal design of pre-training runs for LLMs. However, these considerations neglect inference efficiency. When deploying LLMs at large scale, inference costs and performance are critical aspects. Hence Pope et al. (2023) investigate the inference efficiency of transformer-based LLMs in terms of three criteria: compute, latency, and throughput. More recently Sardana et al. (2024) provided a scaling law analysis of Transformers that extend the pre-training compute optimality consideration (Eq. (1)) to also account for inference compute. This work presents an even more comprehensive analysis in terms of the attainable latency, i.e., time to first token, and the step time during generation. We complement our empirical findings with a quantitative model of a lower bound on time to first token and step time, using the detailed calculation of FLOPs (App. B.3) and MemOps (App. B.4) for both model architectures.
@@
For the prefill stage, the key performance metric is the time to first token (TTFT). Prefill speed is primarily determined by how well the model can maintain a low TTFT while handling large batch sizes and long input sequences. During the generation stage, the key performance metric is the step time, i.e. how long it takes to obtain the next token given the current (potentially batched) sequence. For Transformers, the quadratic complexity of the attention mechanism with respect to the prefill length (App. B.3.3) implies that TTFT is expected to scale quadratically in terms of the prefill length. In terms of step time we expect linear scaling with respect to the prefill length, as each decoding step involves attention over the entire KV cache. For xLSTMs, in contrast, we expect linear scaling of TTFT and step time that is independent of the prefill length.
@@ 4.2 Modeling Inference Runtimes
In our analysis, the inference processes are characterized by the associated number of floating point operations $\text{FLOPs}_{\text{algo}}$ and the number of memory operations $\text{Bytes}_{\text{mem,algo}}$ measured in bytes that are read or written. We provide calculations of these two quantities for xLSTM and for Transformers in Appendix B. Importantly, these calculations capture the difference between xLSTM and Transformers in the dependence of $\text{FLOPs}_{\text{algo}}$ and $\text{Bytes}_{\text{mem,algo}}$ on the context length $T$. Based on these calculated quantities, we model the runtimes associated with the floating point and memory operations as:
@@
where $\alpha_{\text{ eff}{{}}}$ is the effective rate of FLOPs/s, $\beta_{\text{ eff}{{}}}$ is the effective rate of Bytes/s, and $\epsilon$ is a constant overhead when running the inference processes on the GPU. Depending on the model type, model size, prefill length, batch size and inference stage (prefill or generate), either $\tau_{\text{ FLOPs}{{\text{,algo}}}}$ or $\tau_{\text{ mem}{{\text{,algo}}}}$ is the dominant contributor to the runtime. We outline in Appendix C.1 how this is determined based on the roofline model. Using empirical runtime measurements, we then fit one of the two models depending on which one is expected to yield the dominant runtime contribution. Each fit corresponds to a specific model type, size, and inference stage, and is evaluated over varying batch sizes and prefill lengths. As evidenced by the fits to empirical TTFT (App. C.2) and step time measurements (App. C.3), our model provides an accurate description of the observed inference runtimes for both architectures and explains the empirically observed runtimes in Figure 6.
@@ 5 Related Work
The empirical scaling behavior of Deep Learning models w.r.t the size of their model parameters and training data has been actively researched (Hestness et al., 2017; Rosenfeld et al., 2020; Henighan et al., 2020; Alabdulmohsin et al., 2022; Caballero et al., 2023). Such scaling laws have been demonstrated across many tasks and data modalities (Tan and Le, 2019; Ghorbani et al., 2022; Zhai et al., 2022; Abnar et al., 2022; Ardalani et al., 2022; Gao et al., 2023) However, beginning with Kaplan et al. (2020) and Hoffmann et al. (2022), the main objective has been guidance on how to optimally scale Large Language Models with Transformers. Follow-up work investigated the data constrained setting (Muennighoff et al., 2023), the effect of data pruning (Sorscher et al., 2022), extreme token per parameter ratios (Gadre et al., 2024). Furthermore, replication efforts regarding the scaling laws established in Kaplan et al. (2020) and Hoffmann et al. (2022) have been performed in order to reconcile their findings (Besiroglu et al., 2024; Pearce and Song, 2024; Porian et al., 2024). Critical practical considerations such as specific architectures and hyperparameters on the resulting scaling laws have been investigated (McLeish et al., 2025). The recent survey Li et al. (2025) gives a comprehensive overview and give practical guidelines in establishing scaling laws. Scaling laws have also been investigated theoretically, providing justification for the functional forms used in practice (Amari et al., 1992; Amari, 1993; Seung et al., 1992; Amari and Murata, 1993; Cortes et al., 1993; Yarotsky, 2018; Liang et al., 2020; Sharma and Kaplan, 2022; Hutter, 2021; Bahri et al., 2024).